A velocity of magnitude 40m per sec is directed at an angle of 40degree east of north.draw a vector to represent this velocity


1
Expert's answer
2021-01-31T03:41:00-0500

Let's first calculate the horizontal and vertical component's of the velocity vector:


"v_x=vcos\theta=40\ \dfrac{m}{s}\cdot cos40^{\circ}=30.6\ \dfrac{m}{s},""v_y=vsin\theta=40\ \dfrac{m}{s}\cdot sin40^{\circ}=25.7\ \dfrac{m}{s}."

Let's draw the vector (here one division corresponds to 5 m\s ):